# Nonasymptotic Convergence Rates for Plug-and-Play Methods With MMSE Denoisers

### v2

@@ I Background
where $f$ is a data-fidelity term (typically $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$ under Gaussian noise), and $g$ is a regularization term that encodes prior beliefs about the class of admissible solutions. For instance, the $\ell^{1}$-norm ($g(x)\propto\lVert x\rVert_{1}$) promotes sparsity and underlies the theory of compressed sensing [9, 17, 6]. Similarly, total variation (TV), defined as $g(x)\propto\operatorname{TV}(x)\coloneqq\sum_{k}|x_{k}-x_{k-1}|$ favors piecewise-constant solutions and is widely used in image denoising [52, 10, 11, 36, 44]. More recently, much of research in imaging inverse problems has shifted towards learned regularizers [29, 18, 19, 24, 48, 22, 43, 33]. Without regularization, the problem is often ill-posed in the sense of Hadamard [21, 28]: Solutions may be non-unique, unstable, or nonexistent. Since the data-fidelity term is typically well-behaved, the regularizer largely determines the structure of solutions. More broadly, the regularizer largely dictates how optimization algorithms behave, so understanding its structure is crucial.
@@
With appropriately chosen step size $\gamma>0$, the iterates $x_{k}$ converge to a stationary point [32]. However, gradient descent is not well-suited for nonsmooth objectives.
@@
Here $\operatorname{prox}_{\gamma g}$ denotes the proximal operator of $g$ (see Section II), originally introduced by Moreau in [40]. Convergence of PGD has been widely studied in both convex and non-convex settings and can be established under various assumptions on $f$ and $g$ with appropriately chosen step sizes $\gamma>0$ [35, Table 1].
@@ I-B Plug-and-Play Methods
The plug-and-play (PnP) framework [58, 33] extends PGD by replacing the proximal step $\operatorname{prox}_{\gamma g}$ with a generic denoiser $D_{\sigma}$ [7]. This substitution is motivated by the interpretation of $\operatorname{prox}_{\gamma g}$ as the maximum a posteriori (MAP) solution to the denoising problem (2) under prior $p_{X}\propto e^{-\gamma g}$ [14, Eq. (10.13)], the exact problem denoisers are designed to solve. The PnP iteration is given by
@@
and has been shown to yield state-of-the-art results when combined with powerful image denoisers such as BM3D [16] or a deep neural network (DNN) [31, 60]. Often, we use data fidelity term $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$, in which case the PnP iteration becomes
@@
A central challenge in the PnP framework is that while a proximal operator is always tied to a well-defined regularizer, many of the most effective denoising methods are not. This substitution therefore breaks the direct connection to an explicit optimization problem, making theoretical interpretations and convergence claims difficult. Much of the literature imposes strict assumptions on the denoiser to obtain convergence guarantees. Many of the most effective denoisers do not satisfy these assumptions, yet still achieve excellent results (for example [1]).
@@
When $f$ and $g$ are convex with $L$-smooth $f$, it is well known that PGD converges to a fixed point for appropriate step size $\gamma$ [45, Section 4.2]. This result can be proven via the machinery of monotone operator theory, the goal being to show that the composite PGD step (5) constitutes an averaged operator, guaranteeing convergence [15, 13, 3]. [^1] When $g$ is proper, lower-semicontinuous, convex, $\partial g$ is maximally monotone, making $\operatorname{prox}_{\gamma g}=(I+\gamma\partial g)^{-1}$ firmly-nonexpansive, i.e., $1/2$-averaged. The gradient step is also averaged for $L$-smooth $f$, making their composition an averaged operator.
In particular, convergence of PnP-PGD has been shown under tight assumptions on the denoiser, such as averaged/nonexpansive behavior [56] or suitable Lipschitz (e.g., residual-Lipschitz) bounds [53]. Many works impose these assumptions through architectural constraints on the learned denoisers via, e.g., spectral normalization [53] or through explicit training techniques [57, 8]. In [47], the authors propose a technique for learning a maximally monotone operator. Convergence of PnP has also been shown for specific denoisers, such as kernel denoisers [23], denoisers that can be written as the gradient step of a smooth function [30], bounded denoisers [12], and linear denoisers [42].
Instead of treating the denoiser as a black box with certain conditions, a complementary line of work leverages its intrinsic statistical structure. The MMSE estimator under Gaussian noise is a notable example. With the Tweedie/Stein identities [20, 55], the MMSE denoiser can be tied to the score function, endowing it with useful structural properties [38]. Further, [25] and [27] showed that the MMSE denoiser is $C^{\infty}$ and can be written as the proximal map of a regularizer (which is $C^{\infty}$) under various forward and noise models, and derived an explicit formula for this regularizer (28). This progress enabled the derivation of asymptotic convergence for both the ADMM [46] and PGD [59] variants of PnP using an MMSE denoiser. However, a gap in the theory remains: The explicit formula offers little insight into the behavior of the induced regularizer, and there are no nonasymptotic convergence rates tailored to MMSE denoisers.
@@
  • Experiments. In Section IV, we consider the following experimental setups: (i) A one-dimensional synthetic example validating the upper Moreau envelope form of the regularizer implicit in the MMSE denoiser by comparing a learned DNN estimator to its explicit form, (ii) Gaussian deblurring on the MNIST dataset [34], and (iii) computed tomography on the MayoCT dataset [37]. In both imaging tasks we observe the predicted sublinear $\mathcal{O}(1/\sqrt{k}$) convergence in the subgradient residual.
@@ II Preliminaries
In this section, we collect some definitions and results used in the remainder of the paper. First, we recall the definition of the convex conjugate, also known as the Legendre-Fenchel transform or Fenchel conjugate. See for example [51].
@@
Note that the lower Moreau envelope is the usual Moreau envelope, but we make the distinction in this paper since we use both lower and upper Moreau envelopes. For the infimum defining the lower Moreau envelope to be achieved, $f$ must be lower-semicontinuous. An important property of the lower Moreau envelope $M_{\gamma}f$ is that it shares global minimizers with $f$ whenever $f$ is lower-semicontinuous, i.e. $\operatorname*{arg\,min}_{x}f(x)=\operatorname*{arg\,min}_{x}M_{\gamma}f(x).$ See [4] for further background.
@@
Lemma II.5 ([4, Lemma 3]).
@@
and the result follows when we recall that a function equals its biconjugate if and only if it is convex [50, Section 12.2]. ∎
@@
for all $x\in\mathbb{R}^{n}$. When $f$ is a convex, lower-semicontinuous proper function, it is well known that $\operatorname{prox}_{\gamma f}$ and $M_{\gamma}f$ are also related by the following formula due to Moreau [41]:
@@
For a thorough introduction to Tweedie’s formula, we refer the reader to [20]. Note that Tweedie’s Formula makes no assumptions on the prior density $p_{X}$, which may even include discrete atoms.
Theorem II.9 ([27] and [25, Corollary 1]).
@@ III Main Results
When the objective $F$ is smooth, the map $x\mapsto\nabla F(x)$ is continuous. Therefore, measuring convergence via the gradient norm $\|\nabla F(x_{k})\|$ is natural; small gradient norms certify proximity to a stationary point. In the non-smooth, non-convex setting, the gradient may not exist and is generalized to the subgradient [51, Chapter 8].
@@
In non-convex, nonsmooth problems, the stationary residual (41) may oscillate, even while it trends downward. This can depend on the forward operator as well as the data. To account for this behavior, nonasymptotic convergence results are typically stated in terms of the best iterate so far. See for example [2]. It is standard to state that an algorithm reaches an $\varepsilon$-stationary point in the first $k$ iterations if
@@
For our investigation, we prove the following intermediary result, which is a simple improvement over [5, Theorem 5.1]. This result will be used to prove convergence of PGD using an MMSE denoiser.
@@
Convergence of PGD using an MMSE denoiser has been previously established for a range of step sizes $\gamma>0$. In particular, [59] showed that PnP-MMSE converges for any $\gamma\leq 1/L$, but the objective it converges to is not the objective of interest, i.e.,
@@ IV Experiments
Figure 3: MMSE denoiser, convex potential, and implicit MMSE regularizer for a Laplacian prior $p_{X}(x)=e^{-|x|}$, plotted against reference curves. LPN denotes the learned proximal network. This explicates what is shown in [22, Figure 2], but plotted on a wider axis to show the “curling down” behavior of LPNs.
We consider the denoising problem (2) under mixture-of-Gaussian and mixture-of-Laplacian priors with unit Gaussian noise. We train a gradient-output softplus input convex neural network (ICNN) to approximate the MMSE denoiser and calculate its implicit prior via [26, Theorem 3]. We compare this with the implicit regularizer calculated using the explicit upper Moreau envelope formula from Theorem III.2. Figures 1 and 2 compare the calculated $\phi_{\mathrm{MMSE}}$ and the learned estimate, along with reference curves $f_{X}\coloneqq-\log p_{X}$ and $f_{Z}\coloneqq-\log p_{Z}$. The learned and calculated regularizers are nearly identical where the samples are most concentrated.
We observe that the learned implicit regularizer “curls downward” outside the data’s central region. This behavior is intrinsic to ReLU-based neural networks (and their softplus approximations). It is well known that feed-forward networks with piecewise-affine activations yield piecewise-affine mappings [49, 39]. In particular, ReLU is piecewise-affine and is the uniform limit of the softplus family as $\beta\to\infty$. Thus, an ICNN with softplus activations is a smooth approximation of a ReLU-ICNN and is therefore “nearly” piecewise-affine. Outside the support of the training data, the learned convex potential becomes nearly affine, and its gradient (the denoiser) saturates at an approximately constant value. Consequently, the implicit regularizer appears to bend downward outside the central region. We emphasize that this behavior arises as a result of the ReLU-approximating architecture, not from the MMSE loss or our implementation. To illustrate this, in Figure 3 we reproduce the setup of [22, Fig. 2] and simply extend the axis range. The extended view makes the affine-tail/saturation effect apparent.
@@
We next consider PGD for the Gaussian blurring inverse problem (1), where $X$ is sampled from the MNIST data set of handwritten digits [34] (resolution of $28\times 28$). $A$ is a Gaussian blurring operator, computed by convolution with a $3\times 3$ Gaussian kernel with $\sigma^{2}=1$, and measurement noise $\epsilon$ being a zero-mean standard Gaussian with variance $\sigma^{2}=0.04$. The Gaussian kernel has unit mass, giving the blur an operator norm $\lVert A\rVert_{\mathrm{op}}=1$.
@@
We study the CT inverse problem with forward operator $A$ given by the discretized Radon transform, normalized to $\|A\|_{\mathrm{op}}<1$. Training images are random $128\times 128$ patches sampled from the MayoCT dataset [37] (resolution of $512\times 512$).
We test the PGD algorithm using the MMSE denoiser on randomly selected regions of the Shepp–Logan phantom [54], which captures a wide range of structures relevant in computed tomography. We corrupt measurements with zero-mean Gaussian noise to yield input PSNR levels of $20$, $24$, and $28$ dB, ranges commonly studied for low-dose CT (LDCT). See for example [61, Table 6]. After $5$ iterations of PGD we obtain reconstructions in the range of 37373838 dB which we display in Figs. 6 and 7.
@@
We build upon the implementation of the Input Convex Neural Network (ICNN) in [22] to approximate the MMSE denoiser. The authors produced excellent results with softplus activations and a mixture of linear and convolutional layers. [^3] The exact PyTorch implementation is available at https://github.com/ZhenghanFang/learned-proximal-networks The network uses layers of the form
@@
To ensure convexity in $x$, all hidden layers $W_{k}$ are clipped so that they have only non-negative entries. Note that a convolutional layer can be represented as a matrix. For the MayoCT and MNIST experiments, the hidden layers are clipped in this form, not in kernel form. Because the activation is convex and non-decreasing, these constraints make $g_{\theta}$ convex, guaranteeing that its gradient ($\psi_{\theta}$) is a proximal operator [26, Corollary 1]. Note that no restrictions on the skip-connections or biases are necessary.

### v3

@@ I Background
where $f$ is a data-fidelity term (typically $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$ under Gaussian noise), and $g$ is a regularization term that encodes prior beliefs about the class of admissible solutions. For instance, the $\ell^{1}$-norm ($g(x)\propto\lVert x\rVert_{1}$) promotes sparsity and underlies the theory of compressed sensing [candes_robust, donoho_compressed, bora_compressed]. Similarly, total variation (TV), defined as $g(x)\propto\operatorname{TV}(x)\coloneqq\sum_{k}|x_{k}-x_{k-1}|$ favors piecewise-constant solutions and is widely used in image denoising [rudin_nonlinear, chambolle_algorithm, chambolle_first, liu_image, parhi2023sparsity]. More recently, much of research in imaging inverse problems has shifted towards learned regularizers [hertrich2025learning, ducotterd2025learning, ducotterd2025multivariate, goujon2023neural, pourya2025dealing, fang_whats, ongie2020deep, kamilov_plug]. Without regularization, the problem is often ill-posed in the sense of Hadamard [engl_regularization, hadamard_problemes]: Solutions may be non-unique, unstable, or nonexistent. Since the data-fidelity term is typically well-behaved, the regularizer largely determines the structure of solutions. More broadly, the regularizer largely dictates how optimization algorithms behave, so understanding its structure is crucial.
@@
With appropriately chosen step size $\gamma>0$, the iterates $x_{k}$ converge to a stationary point [kakade_dimension]. However, gradient descent is not well-suited for nonsmooth objectives.
@@
Here $\operatorname{prox}_{\gamma g}$ denotes the proximal operator of $g$ (see Section II), originally introduced by Moreau in [moreau_functions]. Convergence of PGD has been widely studied in both convex and non-convex settings and can be established under various assumptions on $f$ and $g$ with appropriately chosen step sizes $\gamma>0$ [li_accelerated, Table 1].
@@ I-B Plug-and-Play Methods
The plug-and-play (PnP) framework [venkatakrishnan_plug, kamilov_plug] extends PGD by replacing the proximal step $\operatorname{prox}_{\gamma g}$ with a generic denoiser $D_{\sigma}$ [boyd_distributed]. This substitution is motivated by the interpretation of $\operatorname{prox}_{\gamma g}$ as the maximum a posteriori (MAP) solution to the denoising problem (2) under prior $p_{X}\propto e^{-\gamma g}$ [combettes_proximal, Eq. (10.13)], the exact problem denoisers are designed to solve. The PnP iteration is given by
@@
and has been shown to yield state-of-the-art results when combined with powerful image denoisers such as BM3D [dabov_image] or a deep neural network (DNN) [jain_natural, zhang_beyond]. Often, we use data fidelity term $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$, in which case the PnP iteration becomes
@@
A central challenge in the PnP framework is that while a proximal operator is always tied to a well-defined regularizer, many of the most effective denoising methods are not. This substitution therefore breaks the direct connection to an explicit optimization problem, making theoretical interpretations and convergence claims difficult. Much of the literature imposes strict assumptions on the denoiser to obtain convergence guarantees. Many of the most effective denoisers do not satisfy these assumptions, yet still achieve excellent results (for example [ahmad_plug]).
@@
When $f$ and $g$ are convex with $L$-smooth $f$, it is well known that PGD converges to a fixed point for appropriate step size $\gamma$ [parikh_prox, Section 4.2]. This result can be proven via the machinery of monotone operator theory, the goal being to show that the composite PGD step (5) constitutes an averaged operator, guaranteeing convergence [combettes_signal, combettes_solving, bauschke_convex]. [^1] When $g$ is proper, lower-semicontinuous, convex, $\partial g$ is maximally monotone, making $\operatorname{prox}_{\gamma g}=(I+\gamma\partial g)^{-1}$ firmly-nonexpansive, i.e., $1/2$-averaged. The gradient step is also averaged for $L$-smooth $f$, making their composition an averaged operator.
In particular, convergence of PnP-PGD has been shown under tight assumptions on the denoiser, such as averaged/nonexpansive behavior [sun_online] or suitable Lipschitz (e.g., residual-Lipschitz) bounds [ryu_plug]. Many works impose these assumptions through architectural constraints on the learned denoisers via, e.g., spectral normalization [ryu_plug] or through explicit training techniques [terris_building, bredies_learning]. In [pesquet_learning], the authors propose a technique for learning a maximally monotone operator. Convergence of PnP has also been shown for specific denoisers, such as kernel denoisers [gavaskar_plug], denoisers that can be written as the gradient step of a smooth function [hurault_proximal], bounded denoisers [chan_plug], and linear denoisers [nair_fixed].
Instead of treating the denoiser as a black box with certain conditions, a complementary line of work leverages its intrinsic statistical structure. The MMSE estimator under Gaussian noise is a notable example. With the Tweedie/Stein identities [efron_tweedies, stein_bound], the MMSE denoiser can be tied to the score function, endowing it with useful structural properties [milanfar_denoising]. Further, [gribonval_reconciling] and [gribonval_should] showed that the MMSE denoiser is $C^{\infty}$ and can be written as the proximal map of a regularizer (which is $C^{\infty}$) under various forward and noise models, and derived an explicit formula for this regularizer (28). This progress enabled the derivation of asymptotic convergence for both the ADMM [park_convergence] and PGD [xu_provable] variants of PnP using an MMSE denoiser. However, a gap in the theory remains: The explicit formula offers little insight into the behavior of the induced regularizer, and there are no nonasymptotic convergence rates tailored to MMSE denoisers.
@@
  • Experiments. In Section IV, we consider the following experimental setups: (i) A one-dimensional synthetic example validating the upper Moreau envelope form of the regularizer implicit in the MMSE denoiser by comparing a learned DNN estimator to its explicit form, (ii) Gaussian deblurring on the MNIST dataset [lecun_mnist], and (iii) computed tomography on the MayoCT dataset [mcollough_low]. In both imaging tasks we observe the predicted sublinear $\mathcal{O}(1/\sqrt{k}$) convergence in the subgradient residual.
@@ II Preliminaries
In this section, we collect some definitions and results used in the remainder of the paper. First, we recall the definition of the convex conjugate, also known as the Legendre-Fenchel transform or Fenchel conjugate. See for example [rockafellar_variational].
@@
Note that the lower Moreau envelope is the usual Moreau envelope, but we make the distinction in this paper since we use both lower and upper Moreau envelopes. For the infimum defining the lower Moreau envelope to be achieved, $f$ must be lower-semicontinuous. An important property of the lower Moreau envelope $M_{\gamma}f$ is that it shares global minimizers with $f$ whenever $f$ is lower-semicontinuous, i.e. $\operatorname*{arg\,min}_{x}f(x)=\operatorname*{arg\,min}_{x}M_{\gamma}f(x).$ See [bernard_lasry] for further background.
@@
Lemma II.5 ([bernard_lasry, Lemma 3]).
@@
and the result follows when we recall that a function equals its biconjugate if and only if it is convex [rockafellar_convex, Section 12.2]. ∎
@@
for all $x\in\mathbb{R}^{n}$. When $f$ is a convex, lower-semicontinuous proper function, it is well known that $\operatorname{prox}_{\gamma f}$ and $M_{\gamma}f$ are also related by the following formula due to Moreau [moreau_proximite]:
@@
For a thorough introduction to Tweedie’s formula, we refer the reader to [efron_tweedies]. Note that Tweedie’s Formula makes no assumptions on the prior density $p_{X}$, which may even include discrete atoms.
Theorem II.9 ([gribonval_should] and [gribonval_reconciling, Corollary 1]).
@@ III Main Results
When the objective $F$ is smooth, the map $x\mapsto\nabla F(x)$ is continuous. Therefore, measuring convergence via the gradient norm $\|\nabla F(x_{k})\|$ is natural; small gradient norms certify proximity to a stationary point. In the non-smooth, non-convex setting, the gradient may not exist and is generalized to the subgradient [rockafellar_variational, Chapter 8].
@@
In non-convex, nonsmooth problems, the stationary residual (41) may oscillate, even while it trends downward. This can depend on the forward operator as well as the data. To account for this behavior, nonasymptotic convergence results are typically stated in terms of the best iterate so far. See for example [attouch_convergence]. It is standard to state that an algorithm reaches an $\varepsilon$-stationary point in the first $k$ iterations if
@@
For our investigation, we prove the following intermediary result, which is a simple improvement over [bohm_variable, Theorem 5.1]. This result will be used to prove convergence of PGD using an MMSE denoiser.
@@
Convergence of PGD using an MMSE denoiser has been previously established for a range of step sizes $\gamma>0$. In particular, [xu_provable] showed that PnP-MMSE converges for any $\gamma\leq 1/L$, but the objective it converges to is not the objective of interest, i.e.,
@@ IV Experiments
Figure 3: MMSE denoiser, convex potential, and implicit MMSE regularizer for a Laplacian prior $p_{X}(x)=e^{-|x|}$, plotted against reference curves. LPN denotes the learned proximal network. This explicates what is shown in [fang_whats, Figure 2], but plotted on a wider axis to show the “curling down” behavior of LPNs.
We consider the denoising problem (2) under mixture-of-Gaussian and mixture-of-Laplacian priors with unit Gaussian noise. We train a gradient-output softplus input convex neural network (ICNN) to approximate the MMSE denoiser and calculate its implicit prior via [gribonval_characterization, Theorem 3]. We compare this with the implicit regularizer calculated using the explicit upper Moreau envelope formula from Theorem III.2. Figures 1 and 2 compare the calculated $\phi_{\mathrm{MMSE}}$ and the learned estimate, along with reference curves $f_{X}\coloneqq-\log p_{X}$ and $f_{Z}\coloneqq-\log p_{Z}$. The learned and calculated regularizers are nearly identical where the samples are most concentrated.
We observe that the learned implicit regularizer “curls downward” outside the central region of the data. This behavior is intrinsic to ReLU-based neural networks (and their softplus approximations). It is well known that feedforward networks with continuous piecewise-affine activations yield continuous piecewise-affine mappings [raghu_expressive, montufar_number]. The ReLU is piecewise-affine and is the uniform limit of the softplus family. Thus, an ICNN with softplus activations is a smooth approximation of a ReLU-ICNN and is therefore nearly piecewise-affine.
Outside the support of the training data, the learned convex potential becomes approximately affine, and its gradient (the denoiser) saturates. To solve for $\phi_{\mathrm{MMSE}}(x)$, we must invert the learned denoiser numerically. This inversion is impossible where the denioser saturates, so we get large values for $\psi_{\theta}^{-1}$ outside the central region. We solve for $\phi_{\mathrm{MMSE}}(x)$ using  [gribonval_characterization, Eq. (8)], so the quadratic $g_{\theta}(\psi_{\theta}^{-1}(x))$ dominates, causing the resulting estimate of the regularizer to curl downward to -$\infty$. We emphasize that this behavior arises as a result of the ReLU-approximating architecture, not from the MMSE loss or our implementation. To illustrate this, in Figure 3 we reproduce the setup of [fang_whats, Fig. 2] and simply extend the axis range. The extended view makes the affine-tail/saturation effect apparent.
@@
We next consider PGD for the Gaussian blurring inverse problem (1), where $X$ is sampled from the MNIST data set of handwritten digits [lecun_mnist] (resolution of $28\times 28$). $A$ is a Gaussian blurring operator, computed by convolution with a $3\times 3$ Gaussian kernel with $\sigma^{2}=1$, and measurement noise $\epsilon$ being a zero-mean standard Gaussian with variance $\sigma^{2}=0.04$. The Gaussian kernel has unit mass, giving the blur an operator norm $\lVert A\rVert_{\mathrm{op}}=1$.
@@
We study the CT inverse problem with forward operator $A$ given by the discretized Radon transform, normalized to $\|A\|_{\mathrm{op}}<1$. Training images are random $128\times 128$ patches sampled from the MayoCT dataset [mcollough_low] (resolution of $512\times 512$).
We test the PGD algorithm using the MMSE denoiser on randomly selected regions of the Shepp–Logan phantom [shepp_fourier], which captures a wide range of structures relevant in computed tomography. We corrupt measurements with zero-mean Gaussian noise to yield input PSNR levels of $20$, $24$, and $28$ dB, ranges commonly studied for low-dose CT (LDCT). See for example [zhao_low, Table 6]. After $5$ iterations of PGD we obtain reconstructions in the range of 37373838 dB which we display in Figs. 6 and 7.
@@
We build upon the implementation of the Input Convex Neural Network (ICNN) in [fang_whats] to approximate the MMSE denoiser. The authors produced excellent results with softplus activations and a mixture of linear and convolutional layers. [^3] The exact PyTorch implementation is available at https://github.com/ZhenghanFang/learned-proximal-networks The network uses layers of the form
@@
To ensure convexity in $x$, all hidden layers $W_{k}$ are clipped so that they have only non-negative entries. Note that a convolutional layer can be represented as a matrix. For the MayoCT and MNIST experiments, the hidden layers are clipped in this form, not in kernel form. Because the activation is convex and non-decreasing, these constraints make $g_{\theta}$ convex, guaranteeing that its gradient ($\psi_{\theta}$) is a proximal operator [gribonval_characterization, Corollary 1]. Note that no restrictions on the skip-connections or biases are necessary.